### v1

# E. Gover11footnotetext: Corresponding author.    D. Veske    M.B. Demirkoz

Abstract
For studies where a reliable neutron source/beam is required and a nuclear reactor is not a viable option (considering their high neutron flux supply, which may not be suitable for research concerning low flux operations), alternative approaches may be sought. This paper presents a comprehensive simulation analysis of 30 MeV ${}^{9}\mathrm{Be}\text{(p,n)}{}^{9}\mathrm{B}$ reaction, which can be utilized as an isotropic neutron source. Due to different underlying physics and transport mechanisms, slight variations in findings between Geant4 and FLUKA may occur, and hence may require one to have a guide at hand to address the differences and interpret the data more accurately before conducting the actual experiment. Through these Monte Carlo simulation toolkits, this work estimates resulting neutron fluence and dose equivalent of prompt gamma decays, as well as how materials such as paraffin and borated polyethylene moderates its energy spectrum. Furthermore, this work also presents an examplar modular irradiation station designed for target-moderator configurations, with the capability of generating thermal neutron fields.
1 Introduction
For studies where a reliable neutron source/beam is required and a nuclear reactor is not a viable option (considering their high neutron flux supply, which may not be suitable for research concerning low flux operations), alternative approaches may be sought. For instance, material hardness studies, clinical neutron radiography, neutron diffraction studies, investigation of neutron induced activation reactions, neutron activation analysis serve as exemplars of the full spectrum of outcomes attainable through this configuration studied in this work. Moreover, SEU (single event upset) tests for electronic boards that will be subject to neutron radiation in upper atmosphere where the dominant source of upsets are neutrons can also be conducted with a novel accelerator-based neutron source. For facilities operating a proton- or deuteron-based accelerators, one method is utilizing the ${}^{9}\mathrm{Be}\text{(p,n)}{}^{9}\mathrm{B}$ reaction branch to construct and orient a neutron beam. For the purpose of optimizing the irradiation parameters and benchmarking prior to tests, simulation codes such as GEANT4[1, 2, 3] and FLUKA[4, 5] shall be deployed. Although both simulation toolkits offer solid and consistent numerical groundwork, due to their dissimilarities in underlying physics models and code structure, there can be slight variations in computational output. This study will examine and compare the variations observed between the toolkits. To allow for an unbiased comparison, all simulation parameters will be kept the same.
2 Target Optimization
Table 1 presents neutron fluences corresponding to various target angular orientations. As $45^{\circ}$ angle results in the highest yield, it was chosen for target orientation.
Table 1: Total fluence response due to different target angles at a given point P(x,0,0) on X axis. FLUKA
Angle[Degrees] Total Response[p/cmq/pr]
$45^{\circ}$ 3.3099E-05
$30^{\circ}$ 2.996E-05
$20^{\circ}$ 2.794E-05
$10^{\circ}$ 2.589E-05
$0^{\circ}$ (Perpendicular target) 2.231E-05
After determining the optimal target angle, we then moved forward to find the optimal incident energy. To do so, spherical neutron counts were obtained for different energies. Using Geant4, the following trend was acquired.
Figure 1: Different incident proton energies were tested to get the neutron count coming out of an inclined target (left). Penetration depth of protons with respective energies (right). Geant4 and FLUKA
Figure 1 suggest the highest neutron production at 26 MeV, but for the sake of reducing the activity resulting from decays of prompt isotopes resulting from the irradiation, 30 MeV was chosen, as higher proton energies indicate lower radioactive contamination in the irradiation chamber.
Table 2: Total neutron counts for different beryllium thickness values (30 MeV p, $45^{\circ}$ target angle). Geant4
Thickness [mm] Counts
2.0 14441
3.0 21244
4.0 24074
5.0 24260
Table 2 shows the total spherical neutron count with respect to different target thicknesses. To overcome the blistering effect of protons, the chosen target thickness should be thinner than the stopping range of protons with 30 MeV energy.
Figure 1 suggests that in order to not deal with blistering of Hydrogen, proton’s range of 5.8 mm should be taken into consideration. Hence, target thickness shall not exceed this value, otherwise gradual buildup of hydrogen gas inside the target volume will occur. The following results were done for a target with thickness 2 mm and 20 mm width, 20 mm length.
3 GEANT4 Configuration and Code Description
3.1 Experimental Setup
Figure 2: Outer geometry showing moderator(green), lead casing and target-beam dump duo.
Figure 3: Inner geometry showing beam dump (light blue) and target(dark blue).
[…]
3.2 Physics
In this work, slight modifications were imposed on radioactivation example that is preconfigured and present in GEANT4 directory. As both this example and our goal handles hadronic cascade reactions, choice of hadron inelastic physics plays an important role as different physics lists can yield different results. For that, QSP_BIC_HP physics list for inelastic and HadronElasticPhysicsHP for elastic scattering have been registered. High precision (HP) addition carries utmost importance as the resulting neutron energy ranges from thermal to fast. HP enables the transport of thermal neutrons (or more generally, neutron energies < 19.5 MeV). G4NDL 3.16 (Neutron Data Library) was set for ENDF/B-VII to retrieve thermal neutron cross section. The complete list of physics list that was registered are as follows:
  • HadronElasticPhysicsHP: Handles elastic scattering of hadrons using high-precision (HP) models, which are particularly important for low-energy neutron interactions.
  • G4HadronPhysicsQGSP_BIC_HP: Manages inelastic scattering of hadrons. The "QGSP" refers to the Quark-Gluon String Model for high-energy interactions, while "BIC" (Binary Cascade) is used for intermediate-energy interactions.
  • G4IonElasticPhysics: Governs elastic scattering processes for ions, such as alpha particles or heavier nuclei, ensuring accurate treatment of nuclear interactions.
  • G4IonPhysicsXS:Handles inelastic interactions of ions, incorporating cross-section data to improve simulation accuracy for processes like spallation or fragmentation
  • G4StoppingPhysics: Deals with stopping mechanisms for short-lived hadrons, including processes like nuclear capture at rest and interactions of low-energy charged particles with matter.
  • GammaNuclearPhysics: Simulates interactions between gamma rays and nuclei, including photonuclear reactions where high-energy gamma photons induce nuclear disintegration.
  • ElectromagneticPhysics: Covers electromagnetic interactions such as ionization, Bremsstrahlung, Compton scattering, and pair production, ensuring accurate modeling of charged particle and photon interactions
  • G4DecayPhysics: Implements decay processes for unstable particles, including weak decays of elementary particles like muons and mesons.
  • RadioactiveDecayPhysics: Models radioactive decay of nuclei, including alpha, beta, and gamma emissions, as well as spontaneous fission. This will give better sense for dose calculations around the target.
  • G4NeutronTrackingCut: Applies tracking cuts to low-energy neutrons to improve simulation efficiency by removing very slow neutrons that would not contribute significantly to the results.
3.3 Fluence Estimation Principle
To interpret and analyze the wide spectrum of neutron energy, logarithmic binning was essential. The binning in Geant4 is done manually in post calculation unlike the user routines feature in FLUKA. Results were acquired in counts, hence the redistribution of neutron counts to their corresponding energy bin intervals was done by a ROOT macro file. The logarithmic energy binning was applied according to[6]:
$E_{i}=E_{\min}\times\left(\frac{E_{\max}}{E_{\min}}\right)^{\frac{i}{N}}$ (3.1)
This ensured both FLUKA and Geant4 gave results in the same dimensions and physical quantities.
4 FLUKA Configuration and Code Description
4.1 Experimental Setup
Figure 4: Constructed geometry using FLAIR and GeoViewer, graphical interface of FLUKA. [7].
In Figure 4, a 3D section view of the proposed experiment is shown. Target composition is 100% pure beryllium. The isotropic fraction of beryllium is also assumed to be 100% ${}^{9}\mathrm{Be}$. The choice of target was decided according to gamma/neutron ratio and thermal conductivity of preferred target material. ${}^{7}\text{Li}$ yields better neutron production results.[8, 9] but has lower thermal conductivity and higher gamma yield. With respect to beam direction, the target is rotated with 45-degree angle. This was found to be the optimal positioning after various runs with different angles. This results in slight diminishment in back scattering of neutrons. resulting neutron flux was registered just before the moderation block, where only neutron steps that enter the boundary were counted. Immediately downstream of the target, an aluminum beam dump is strategically positioned. This component fulfills two critical functions. It effectively intercepts and absorbs the incident proton beam, halting its trajectory, and the target material exhibits a thermal conductivity of 210 W/mK[10], rendering it amenable to efficient water cooling. The aluminum beam dump facilitates this cooling process by acting as a heat sink, dissipating the energy deposited in the target to prevent melting (beryllium is highly corrosive and toxic). This implies proton-aluminum reaction will also contribute to neutron fluence. For attenuation of gamma rays, 5 cm thick lead shell encapsulates the whole setup. To accurately simulate the radiation environment to correctly assess neutron fluence and radiation dose within a cyclotron research and development setting, the construction of the irradiation chamber necessitates the incorporation of substantial shielding. In alignment with the design principles observed at the TENMAK-NUKEN Proton Accelerator Facility in Turkey, a concrete wall thickness of 2.5 meters is deemed appropriate for this purpose. In FLUKA, this material is defined as Portland cement, which is present in material library by default.
Figure 5: Top view of the irradiation room. The outer gray box represents concrete shielding.
4.2 Physics
All runs were done with FLUKA version 4-3.3. For base level predefined physics settings, “PRECISION” option was selected. This option ensures low energy neutrons are properly transported (below 20 MeV). As for additional physics, COALESCE card was activated to ensure light cluster formation takes place, which is also included in FLUKA’s PEANUT model (Pre-Equilibrium and Evaporation Nuclear Model). This model is particularly useful for nuclear inelastic interactions for neutron energies above 20 MeV. Below this threshold, nuclear elastic scatterings are based on Ranft model (see [11]) and FLUKA follows group-wise approach for treatment of neutrons. This means at each energy group, reaction cross-sections are taken from evaluated nuclear databases such as ENDF, JEFF, JENDL are averaged over neutron fluence. Cross-sections calculations are hence concluded by equation (4.1)
[…]
Due to the projectile particle being a proton beam having energy in the order of few MeV, the coalescence process could result in the formation of light nuclei like, ${}^{2}\mathrm{H},{}^{3}\mathrm{H},{}^{3}\mathrm{He},{}^{4}\mathrm{He}$ which can indirectly contribute to neutron and gamma yield via further interactions[12]. To count neutrons more accurately, decays to neutrons shall also be taken into consideration, hence DECAYS card was also activated. Another neutron source is the nuclear evaporation process. Neutron emission from excited nuclei was activated using EVAPORAT card. All calculations were done using built-in scoring by utilization of various estimators.
4.3 Fluence Estimation Principle
Flux calculations in FLUKA inside volumes were handled by USRTRACK, where it estimates particle fluence by dividing the total track length of the responding particle by the volume.
$\dot{\Phi}(v)dt=n(v)vdt=\frac{dN(v)}{dV}\frac{dl(v)}{dt}dt=\lim_{\Delta V\rightarrow 0}\frac{\sum_{i}l_{i}(v)}{\Delta V}$ (4.2)
Which is in simple terms, just
Total Track Length of NeutronsTotal $\frac{\textit{Total Track Length of Neutrons}}{\textit{Total Volume}}=\phi[\text{particles }cm^{-2}]$ (4.3)
As the intended purpose of this proposed setup is for experiments concerning neutron irradiation without regard for net directional flow of neutrons, results will be calculated in terms of zeroth order of neutron angular flux $\Psi(r,\hat{\Omega},E,t)$ where the neutron flux can be found as [13]:
[…]
This gives total neutron population at a given point, regardless of their direction. The current, which is the first order moment of angular flux can be found as:
[…]
which is obtained by weighting $\Psi$ with the direction $\hat{\Omega}$. This accounts for directional flow of neutrons. It can be seen that the fluence is independent of surface orientation, while current is not. We will be interested in $\phi(r,E,t)$.
5 FLUKA Results
Ideal target orientation was determined by analyzing double differential fluence of neutrons using USRBDX detector card. Due to the quasi-logarithmic group energy structure of neutrons, logarithmic energy binning was imposed with 200 energy bins. For the solid angle binning, we set up 15 angular bins, starting from 0 to $\pi/2$. This angle is with respect to surface normal (i.e. target). As the target itself is positioned 45-degree angle with respect to origin, another 45-degree angle with respect to surface normal is equivalent to 90 degrees angle with respect to beam direction(i.e. polar angle $\theta$ where z axis is the beam direction).
Next, we calculated the angular distribution of neutrons with respect to azimuthal angle. As illustrated in Figure 6, neutron scattering events within a target material induce a deviation in the neutron flux trajectory, shifting it towards a more orthogonal direction. The figure on the right shows the high tendency for higher angles. For neutrons of higher energies, scattering emerges as the primary mechanism governing the channeling (or collimation) of the neutron beam. At lower energies, we observe the exact opposite. Lower energy neutrons exhibit a greater tendency to travel along paths parallel to the coordinate axes, hence motion along the directions defined by the orthogonal coordinate axes is more prevalent at lower neutron energies.
Figure 6: Azimuthal ($\phi$) distribution of resulting neutrons.
In order to compare how neutron output fluence behaves under various materials and moderation stages, initial resulting neutron count from proton-beryllium reaction was calculated using USRTRACK estimator. A USRTRACK card was defined for target, which has dimensions of 3 cm x 3 cm with 0.2 cm thickness. This results in a volume of 1.8 $\text{cm}^{3}$ which is used by USRTRACK to normalize fluence (integrated over solid angle $\Omega$). Figure 7 shows the resulting energy spectrum and fluence of neutrons inside the target.
Figure 7: Single differential neutron fluence inside beryllium target, normalized with volume.
Distance from the target is virtually 0, hence this can be treated as the total number of neutrons resulting from beryllium only. In order to properly evaluate the neutron fluence data, a lethargy plot was also given in Figure 8. There is a considerable amount of neutrons present at significantly lower energies, extending to $\sim 10^{-8}$ MeV.
Figure 8: Lethargy plot of neutron fluence inside beryllium target.
The general neutron fluence across the irradiation chamber was calculated using USRBIN. While little neutron leakage from the chamber is expected, with neutrons penetrating the surrounding walls to a certain depth, the concrete shielding is adequate to effectively attenuate neutron flux, as it can be seen in Figure 9. Consequently, no neutron readings are anticipated beyond the confines of concrete shielding. Any neutron detection outside the shielding should therefore yield a background level reading.
Figure 9: Neutron fluence across the room.
To calculate the dose inside the chamber, where neutron and gamma radiation contribute individually, USRBIN was utilized as well. Figure 10 shows the gamma equivalent dose inside the chamber. The results were given in the units of pSv/primary. The dose extends the neutron irradiation port, which has an opening that lacks lead shielding. As the outer lead shell has an opening after borated polyethylene moderator, gamma dose has higher value on that corresponding plane. Same as neutrons, little to no gamma readings are anticipated beyond concrete walls.
Figure 10: Gamma and neutron equivalent doses inside the room.
6 Geant4 Results
By registering the neutrons and their energy inside target volume, the resulting energy spectrum was plotted. Figure 11 shows neutron flux per primary, normalized with volume. Figure 11 shows the resulting spectral fluence of neutrons inside the target. At lower energies (< 1 MeV), FLUKA and Geant4 agrees on neutron spectral fluence. At higher energies, they tend to deviate from each other. A fluctuation between $\sim 10^{-8}$ and $\sim 10^{-2}$ MeV is visible in both Figure 11 and Figure 8.
Figure 11: Lethargy plot (left) and single differential linear plot (right) of neutron spectral fluence inside beryllium target. Geant4
Figure 12 shows the moderation done by polyethylene block of 20 cm. The observed increase in thermal neutron concentration, as evidenced by the heightened intensity of closer peaks in lower energy bands, suggests a moderating effect induced by the presence of polyethylene.
Figure 12: Moderated neutron flux. Geant4
A thin spherical shell was used to encapsulate beryllium target to register the angular distribution of outgoing neutrons. A peak in neutron count around $\theta$ $\sim 70$ was observed. The probability density function of $\Theta$ depicted in Figure 13 (right) demonstrates a leftward spectral shift within the visible range. This shift, interpreted as a consequence of the initial target orientation, indicates a contribution of 45 degrees to the observed beam orientation.
Figure 13: Spherical angular distribution of neutrons emerging from beryllium target. Geant4.
Figure 14 shows overall gamma flux $E\ \Phi(E)/dE$
Figure 14: Resulting gamma flux spectrum. Y-axis is shown in logarithmic (left) and linear (right) scale.
7 Neutron Beam Distribution Characteristics
Figure 15 shows the beam profile projected onto YZ plane. Constructed neutron beam was found to follow a Gaussian profile. To determine the fitted 2D Gaussian of the neutron beam , we used[14]
$f(z,y)=A\exp\left(-\left(\frac{\left(z-z_{0}\right)^{2}}{2\sigma_{Z}^{2}}+\frac{\left(y-y_{0}\right)^{2}}{2\sigma_{Y}^{2}}\right)\right)$ (7.1)
where the fitted parameters were extracted from ROOT file. The calculated Gaussian parameters were as follows: A=749.252, $y_{0}$=0.009, $z_{0}$=2.942 with standard deviations $\sigma_{y}$=14.874, $\sigma_{z}$=14.220. For comparison, the FWHM of incoming proton beam was 1.5 cm with $\sigma=0.636$. Figure 16 and 17 shows the LEGO plot of beam profile in 2D and 3D respectively. The neutron beam projection spans a large area, which extends to $\sim 100\text{ cm}$ in each direction. To better collimate the beam and obtain better focused neutron distribution, novel collimator designs should be implemented.
Figure 15: Neutron Fluence Projection and Fitted 2D Gaussian Distribution
Figure 16: Gaussian beam profile given in probability distribution across YZ plane (COLZ representation).
Figure 17: Gaussian beam profile given in probability distribution across YZ plane (LEGO representation).
8 Conclusion
This work presented the preliminary trial run of a proposed proton-beryllium irradiation experiment. Different nuclear simulation toolkits may present different results, hence this study was conducted to address the differences between FLUKA and Geant4. All simulation results in this work should be cross-checked with experimental beryllium irradiation data from EXFOR[15]. The numerical data obtained from FLUKA using the PRECISION physics as default setting is in more parallel with Geant4 when the physics list specified in Section 3.2 is registered.
Beryllium is a compelling target material for accelerator-driven neutron sources. The significant flexibility in tailoring the neutron energy spectrum (which can be done by varying the incident particle energy in the accelerator and the moderators) enables the feasibility of a wide range of neutron-based studies. Moreover, the isotropic nature of the produced neutron distribution is dependent on both the target orientation, and the kinetic energy of the incoming projectile. This angular aspect provides an additional degree of freedom for shaping the neutron beam. The low atomic number of beryllium (Z=4) and relatively low neutron binding energy (2.4 MeV) enables efficient neutron production through reactions such as ${}^{9}\mathrm{Be}(p,n)^{9}\mathrm{B}$ and ${}^{9}\mathrm{Be}(d,n)^{10}\mathrm{B}$. Its relatively high melting point and thermal conductivity makes it convenient to implement liquid water as primary cooling mechanism, which would be an efficient and cheap alternative to cryogenic cooling. However, beryllium should be handled with caution, as it is highly toxic, volatile under high-intensity irradiation and requires strict safety protocols.
Acknowledgments
The author expresses gratitude to Dr. Doğa Veske and Prof. Dr. M. Bilge Demirköz for their valuable support throughout this work.